# SyncDiff: Synchronized Motion Diffusion for Multi-Body Human-Object Interaction Synthesis

### v4

@@ C.3 A Brief Introduction to BPS Algorithm
Its working principle is as follows: First, a large enough sphere is chosen such that when its center coincides with any object’s centroid, it can fully contain the object. In practice, we choose radius $r=1\text{m}$. Then, $1024$ points are randomly sampled from the sphere’s volume. BPS representation is computed by calculating the difference from each sampled point to the nearest point on the object’s surface. This results in a vector of size $\mathbb{R}^{1024\times 3}$.
@@ C.4 Details for Ablation Studies
In our experiments in Section  4, there are three categories of ablation studies: “w/o decompose”, “w/o $\mathcal{L}_{\text{align}}$”, and “w/o exp sync”.
@@ C.5 Pseudocode for CSR, CRR, and CSIoU
12:procedure CSIoU($o$, $h$, $o^{{}^{\prime}}$, $h^{{}^{\prime}}$)
13:     $\text{CSIoU}\leftarrow 0$
@@ D.1 Dataset Statistics
The sizes of each dataset split are shown in Table  12.
@@ D.2 Hyperparameters in Model Architecture
Action/object Label Feature Extraction Branch. Action/object label features are first encoded by pretrained CLIP [64], and then pass through a 2-layer MLP. Specific parameters are shown in Table  13.
@@
Raw Feature Space Dimension $512$
MLP Architecture Linear($512$, $512$)
@@
Object Geometry Feature Extraction Branch. Object geometry features are first encoded by pretrained BPS [59], and then pass through a 2-layer MLP. Specific parameters are shown in Table  14.
@@
Raw Feature Space Dimension $1024\times 3$
MLP Architecture Linear($1024\times 3$, $512$)
@@
Noise Timestep Embedding Module. Detailed architecture is shown in Table  15.
@@
Timestep Embedder SinusoidalPosEmbedding($64$)
Linear($64$, $256$)
@@
Transformer Encoder-Decoder. Concatenate action label features, object label features (for $m$ rigid objects), object geometry features (for $m$ rigid objects), and the shape parameters $\beta$ of $n$ articulated skeletons to form a condition vector, whose shape is $\mathbb{R}^{128\times(2m+1)+10n}$, as is mentioned in Section  3.4. After replicating and concatenating with padding mask, $x_{\text{dc}}$ and $x_{\text{F}}$, the dimension becomes $\mathbb{R}^{N\times\left(128\texttimes(2m+1)+10n+7m+3Dn+7m(m-1)+3Dmn+1\right)}=\mathbb{R}^{N\times\left(128\texttimes(2m+1)+10n+D_{\text{sum}}+1\right)}=\mathbb{R}^{N\times\mathcal{C}}$. The architecture of the transformer encoder-decoder is shown in Table  16.
@@
Encoder Conv1D($\mathcal{C}$, $512$)
Latent Transformer $4$-layer, $8$-head, $1024$-dim
Decoder Linear($512$, $D_{\text{sum}}$)
@@
Note that after $x_{\text{dc}}$ and $x_{\text{F}}$ (concatenating with the condition vector) passes through the encoder, their shapes becomes $N\times 512$. Concatenating them together results in a latent vector of shape $N\times 1024$, which can be affiliated by the noise embedding, whose dimension is $1024$.
@@ D.3 Training and Inference Hyperparameters
Other important hyperparameters for training and inference process are shown in Table  17.
@@
Epoch $250$k $140$k $280$k $100$k $100$k
$\lambda_{\text{DC}}$ $1$ $1$ $1$ $1$ $1$
$\lambda_{\text{AC}}$ $2.5$ $0.8$ $0.8$ $0.3$ $0.8$
$\lambda_{\text{norm}}$ $0.1$ $0.1$ $0.1$ $0.1$ $0.1$
$\lambda_{\text{align}}$ $0.3$ $0.3$ $0.3$ $0.15$ $0.15$
$\lambda_{\text{exp}}$ $0.3$ $0.3$ $0.3$ $0.3$ $0.3$
@@ D.4 Time and Space Cost, Hardware Configurations
Time and Space Cost for Training and Inference. The training time, average inference time per sample, and GPU memory usage during training are detailed in Table  18.
@@
Training $20.7$h $9.5$h $40.1$h $7.9$h $3.7$h
Inference $7.7$s $6.5$s $7.3$s $6.6$s $4.2$s
Memory $6.11$G $5.07$G $8.59$G $4.93$G $3.67$G
@@
Time and Space Cost for Mesh Reconstruction. Time and space cost of performing mesh reconstruction for a motion sequence of $N=200$ frames are shown in Table  19.
@@ E.2 Discussions and Some Extensions
This section aims to illustrate some points where readers may have questions or misconceptions. For scenarios where SyncDiff can be extended, we will also offer intuitive methods for expansion.
@@ 1. Does SyncDiff lose the flexibility of “one model for all datasets”?
2. Why doesn’t SyncDiff use the 6 DoF rotation representation that is widely employed in the recent motion synthesis methods?
To represent rotation $R\in SO(3)$ of an object, the 6 DoF representation proposed in  [126] concatenates the first two columns of the $3\times 3$ rotation matrix into a 6-dimensional vector $r$ as the representation. After the model generates the predicted result $\hat{r}$, it is split into two column vectors, normalized, and subjected to Gram-Schmidt orthogonalization, and then completed into the rotation matrix $\hat{R}$ in $SO(3)$.
In our method, since the relative representations need to be solved frequently, if we use such 6 DoF representation, it is necessary to complete them into matrices for operations such as inversion and composition, which significantly increases computational cost. By contrast, quaternions are more convenient for normalization and calculation.
@@
As is stated in Section  2.2, several prominent works inject external priors or constraints into synthesized results by performing linear fusions, imputations or inpaintings during the inference process of diffusion models. The term $-\nabla_{\textbf{x}_{0}}G_{z}(P_{x}^{z}\textbf{x})$ in GMD and the gradient of negative log-likelihood (actually equivalent to the score functions) in SyncDiff (Refer to Section  A.2) both provide the direction of denoising (for both), adhering to constraints (for GMD), or synchronization (for SyncDiff).
@@ 4. Can the articulated skeletons in SyncDiff be other body parts besides hands and the full body?
In Section  3.1, our definition of articulated skeletons is “a set of joints whose motion can be reconstructed based on joint information and shape parameter $\beta$”. To extend SyncDiff to HOI motion synthesis of other body parts, two issues need to be addressed:

### v5

@@ C.3 A Brief Introduction to BPS Algorithm
Its working principle is as follows: First, a large enough sphere is chosen such that when its center coincides with any object’s centroid, it can fully contain the object. In practice, we choose radius $r=1\text{m}$. Then, 102410241024 points are randomly sampled from the sphere’s volume. BPS representation is computed by calculating the difference from each sampled point to the nearest point on the object’s surface. This results in a vector of size $\mathbb{R}^{1024\times 3}$.
@@ C.4 Details for Ablation Studies
In our experiments in Section 4, there are three categories of ablation studies: “w/o decompose”, “w/o $\mathcal{L}_{\text{align}}$”, and “w/o exp sync”.
@@ C.5 Pseudocode for CSR, CRR, and CSIoU
[…]
@@ D.1 Dataset Statistics
The sizes of each dataset split are shown in Table  14.
@@ D.2 Hyperparameters in Model Architecture
Action/object Label Feature Extraction Branch. Action/object label features are first encoded by pretrained CLIP [64], and then pass through a 2-layer MLP. Specific parameters are shown in Table  15.
@@
Raw Feature Space Dimension 512512512
MLP Architecture Linear(512512512, 512512512)
@@
Object Geometry Feature Extraction Branch. Object geometry features are first encoded by pretrained BPS [59], and then pass through a 2-layer MLP. Specific parameters are shown in Table  16.
@@
Raw Feature Space Dimension 1024×31024$\times$ 31024 × 3
MLP Architecture Linear(1024×31024$\times$ 31024 × 3, 512512512)
@@
Noise Timestep Embedding Module. Detailed architecture is shown in Table  17.
@@
Timestep Embedder SinusoidalPosEmbedding(646464)
Linear(646464, 256256256)
@@
Transformer Encoder-Decoder. Concatenate action label features, object label features (for $m$ rigid objects), object geometry features (for $m$ rigid objects), and the shape parameters $\beta$ of $n$ articulated skeletons to form a condition vector, whose shape is $\mathbb{R}^{128\times(2m+1)+10n}$, as is mentioned in  3.4. After replicating and concatenating with padding mask, $x_{\text{dc}}$ and $x_{\text{F}}$, the dimension becomes $\mathbb{R}^{N\times\left(128\texttimes(2m+1)+10n+7m+3Dn+7m(m-1)+3Dmn+1\right)}=\mathbb{R}^{N\times\left(128\texttimes(2m+1)+10n+D_{\text{sum}}+1\right)}=\mathbb{R}^{N\times\mathcal{C}}$. The architecture of the transformer encoder-decoder is shown in Table  18.
@@
Encoder $\mathcal{C}$, 512512512)
Latent Transformer 444-layer, 888-head, 102410241024-dim
Decoder Linear(512512512, $D_{\text{sum}}$)
@@
Note that after $x_{\text{dc}}$ and $x_{\text{F}}$ (concatenating with the condition vector) passes through the encoder, their shapes becomes $N\times 512$. Concatenating them together results in a latent vector of shape $N\times 1024$, which can be affiliated by the noise embedding, whose dimension is 102410241024.
@@ D.3 Training and Inference Hyperparameters
Other important hyperparameters for training and inference process are shown in Table  19.
@@
Epoch 250250250k 140140140k 280280280k 100100100k 100100100k
$\lambda_{\text{DC}}$ 111 111 111 111 111
$\lambda_{\text{AC}}$ 2.52.52.5 0.80.80.8 0.80.80.8 0.30.30.3 0.80.80.8
$\lambda_{\text{norm}}$ 0.10.10.1 0.10.10.1 0.10.10.1 0.10.10.1 0.10.10.1
$\lambda_{\text{align}}$ 0.30.30.3 0.30.30.3 0.30.30.3 0.150.150.15 0.150.150.15
$\lambda_{\text{exp}}$ 0.30.30.3 0.30.30.3 0.30.30.3 0.30.30.3 0.30.30.3
@@ D.4 Time and Space Cost, Hardware Configurations
Time and Space Cost for Training and Inference. The training time, average inference time per sample, and GPU memory usage during training are detailed in Table  20.
@@
Training 20.720.720.7h 9.59.59.5h 40.140.140.1h 7.97.97.9h 3.73.73.7h
Inference 7.77.77.7s 6.56.56.5s 7.37.37.3s 6.66.66.6s 4.24.24.2s
Memory 6.116.116.11G 5.075.075.07G 8.598.598.59G 4.934.934.93G 3.673.673.67G
@@
Time and Space Cost for Mesh Reconstruction. Time and space cost of performing mesh reconstruction for a motion sequence of $N=200$ frames are shown in Table  21.
@@ E.2 Discussions and Some Extensions
This section aims to enumerate some points where readers may have questions or misconceptions, providing detailed explanations. For scenarios where SyncDiff can be extended, we will also offer intuitive methods for expansion.
@@ 1. Does SyncDiff lose the flexibility of “one model for all datasets”?
2. Why doesn’t SyncDiff use the 6 DoF rotation representation that is widely employed in the current motion synthesis methods?
To represent the rotation $R\in SO(3)$ of an object, the 6 DoF representation proposed in  [126] concatenates the first two columns of the $\times 3$ 3 rotation matrix into a 6-dimensional vector $r$ as the representation. After the model generates the predicted result $\hat{r}$, it is split into two column vectors, normalized, and subjected to Gram-Schmidt orthogonalization, and then completed into the rotation matrix $\hat{R}$ in $SO(3)$.
In our method, since the relative representations need to be solved frequently, if we use such 6 DoF representation, it is necessary to complete them into matrices for operations such as inversion and composition, which significantly increases the computational cost. Therefore, we adopt quaternions, which are more convenient for normalization and calculation.
@@
As is stated in Section 2.2, several prominent works inject external priors or constraints into synthesized results by performing linear fusions, imputations or inpaintings during the inference process of diffusion models. The term $-\nabla_{\textbf{x}_{0}}G_{z}(P_{x}^{z}\textbf{x})$ in GMD and the gradient of negative log-likelihood (actually equivalent to the score functions) in SyncDiff (Refer to Section A.2) both provide the direction of denoising (for both), adhering to constraints (for GMD), or synchronization (for SyncDiff).
@@ 4. Can the articulated skeletons in SyncDiff be other body parts besides hands and the full body?
In Section 3.1, our definition of articulated skeletons is “a set of joints whose motion can be reconstructed based on joint information and shape parameter $\beta$”. To extend SyncDiff to HOI motion synthesis of other body parts, two issues need to be addressed:
@@
6. Is the metric RA (Recognition Accuracy) fair? Since the discriminator is trained on the combination of train, val, and test folders, is there any risk of overfitting?
Training the classifier solely on the training split would introduce a more serious issue: motion synthesis models overfitting to training distribution would achieve inflated RA scores. In fact, experiments reveal that this approach causes ground-truth test splits to underperform some baselines. This is because the distributional difference between the ground-truth test splits and train splits is greater than that between the distribution fitted by the generative model on train split and the train split itself.
The fundamental challenge is that any classifier—regardless of its training data—will inherently favor in-domain motions. While a huge and more diverse dataset could mitigate this bias, it is currently impractical due to the acquisition cost of mocap data. A motion recognition foundation model might help mitigate this issue. However, currently there are no such foundation models capable of performing high-quality action recognition based solely on trajectories rather than RGB video inputs.
Besides, the relative rankings produced by our RA metric (Tables 1-4 in main text, Table  4) align with perceptual judgments in user studies (Figure  9), supporting its validity.